An area of 52 square feet is to be enclosed by two squares whose sides are in the ratio of 2:3. Find the sides of the two squares.
Answer by checkley77(12844)

(2X)^2+(3X)^2=52
4X^2+9X^2=52
13X^2=52
X^2=52/13
X^2=4
X=2 ANS FOR THE SIDES OF THE SMALLER SQUARE.
PROOF:
(2*2)^2+(3*2)^2=52
4^2+6^2=52
16+35=52
52=52